OpenStudy (anonymous):

ok, correct me if I'm wrong. If my problem, is,say, (this is totaly not the problem on my homework) the distance between (0, -2), (-5, -1), the answer would be about 7.1, right?

jimthompson5910 (jim_thompson5910):

that's a bit too big

OpenStudy (anonymous):

So, what would you're answer be?

jimthompson5910 (jim_thompson5910):

focus on the x coordinates for a moment

jimthompson5910 (jim_thompson5910):

what is the distance from x = 0 to x = -5 ?

OpenStudy (anonymous):

5

jimthompson5910 (jim_thompson5910):

now focus on the y coordinates: what is the distance from y = -2 to y = -1 ?

OpenStudy (anonymous):

1

jimthompson5910 (jim_thompson5910):

so what we have is a right triangle 

jimthompson5910 (jim_thompson5910):

How do we find z?

OpenStudy (anonymous):

1^2=5^2

jimthompson5910 (jim_thompson5910):

I think you meant to say 1^2 + 5^2 = z^2

OpenStudy (anonymous):

yeah

jimthompson5910 (jim_thompson5910):

z = ??

OpenStudy (anonymous):

the square rout of 26, witch I think is...

jimthompson5910 (jim_thompson5910):

equal to what approximate value?

OpenStudy (anonymous):

5.1?

OpenStudy (anonymous):

ish

OpenStudy (anonymous):

OMG you're my new besty thanks so much!

jimthompson5910 (jim_thompson5910):

here is more practice with the distance formula http://www.purplemath.com/modules/distform.htm